OpenStudy (anonymous):

how do i solve f(x)=2t=t squared

OpenStudy (unklerhaukus):

t=0

OpenStudy (unklerhaukus):

or \[2=t\]

OpenStudy (unklerhaukus):

i assuming you intended to but two equal signs

OpenStudy (anonymous):

yes i got that answer but my book say - infinity to infinity why is that instead of 0 to infinity

OpenStudy (unklerhaukus):

negative infinity does not work

OpenStudy (anonymous):

why does it not work?/

OpenStudy (anonymous):

problem says find the domain of f(t)=2t+t squared i'm confused

OpenStudy (unklerhaukus):

well \[f(x)=2t=t^2 \] \[f(-∞)=2\times-∞=-∞\] however\[(-∞)^2=∞\] (negative times negative is positive

OpenStudy (anonymous):

I see now... so that makes the domain -infinity to infinity